OpenStudy (anonymous):

If you roll two dice, what is the theoretical probability of getting an 8 or above?

OpenStudy (anonymous):

No sorry

OpenStudy (anonymous):

i will give a medal if you can do it

OpenStudy (anonymous):

We check the possibilities when the sum is less than 8, which is when 1+1,1+2,2+1,2+2,2+3,3+2,3+3,3+4,4+3 hence total 9 points. Total number of points in event space is 6^2=36. the probability that the sum is less than 8 is 9/36=1/4. Required probability=1-1/4=3/4

OpenStudy (anonymous):

Check the solution

OpenStudy (anonymous):

3/4 is the answer?

OpenStudy (anonymous):

Yes